OpenStudy (anonymous):

if log(x-3)+log(x+4) - logx = log 5 then what is the solution set

OpenStudy (anonymous):

log[(x - 3)(x + 4)] - logx = log5 log(x^2 + x - 12) - logx = log5 log[(x^2 + x - 12)/x] = log 5 (x^2 + x - 12)/x = 5 x^2 + x - 12 = 5x x ^2 - 4x - 12 = 0 (x + 2)(x - 6) = 0 x = -2, 6 Eliminate the -2 x = 6

OpenStudy (anonymous):

did u factor in the begining

OpenStudy (anonymous):

No, I multiplied with FOIL, which is the opposite of factoring. Try not to mix them up.

OpenStudy (anonymous):

o ok thats wut i meeant

OpenStudy (anonymous):

thx